OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (hba):

Part 1) Just plug in g(x)=6x in place of x in f(x) f(x)=5x-26 f(g(x)) or f(6x)=

OpenStudy (hba):

@Mathhater94 Trying solving it :)

OpenStudy (anonymous):

let me work it out on my paper and i will let ya know what i get

OpenStudy (anonymous):

i got -26 as the answer

OpenStudy (hba):

Sure :)

OpenStudy (hba):

What ? Show me your working ? You have to plug in 6x in place of x in f(x)

OpenStudy (anonymous):

i thought the problem would be set up as f(6x)=5x-26?? did i set it up wrong?

OpenStudy (hba):

Yeah you set it up right Now just put 6x in place of x :)

OpenStudy (anonymous):

oh oh ok i see what i did let me re do lol sorry

OpenStudy (hba):

No worries c:

OpenStudy (anonymous):

f(6x)=5x-26 f(6x)=5(6x)-26 f(6x)=30x-26.... if this is right so far where do i go from here and if it isnt right can you show me where i went wrong??

OpenStudy (hba):

Looks good to me.

OpenStudy (anonymous):

so 30x-26 is the answer right?

OpenStudy (hba):

Maybe.

OpenStudy (hba):

Anything else i can help you with ?

OpenStudy (hba):

@Mathhater94

OpenStudy (anonymous):

no i think i understand now thank you